Automotive antifreeze consists of ethylene glycol, CH

2

(OH)CH

2

(OH), a nonvolatile nonelectrolyte. Calculate the


boiling point and freezing point of a 25.0 wt% solution of ethylene glycol in water.



1
Expert's answer
2022-02-07T22:34:02-0500

FREEZING POINT =-10.3

BOILING POINT =102.8


Explanation:


Freezing point of solution =  - 10.3°C


Boiling point of solution  = 102.8°C


Explanation:


This is about colligative properties:


Freezing point depression → ΔT =  Kf . m


ΔT =  Freezing point of pure solvent - Freezing point of solution


Boiling point elevation → ΔT = Kb . m


ΔT =  Boiling point of solution - Boiling point of pure solvent


Let's determine m which means molality (moles of solute in 1 kg of solvent)


Solute: Ethylene glycol → Mass = 25 g


Moles = Mass / Molar mass → 25 g / 60g/mol = 0.416 moles


Solvent: Water


25 % means, 25 g of solute in 100 g of solution


Therefore, the mass of water is 75 g (100 -25) . (Solution = Solute + Solvent)


We convert the mass from g to kg → 75 g . 1kg /1000g = 0.075 kg


Molality (mol/kg) = 0.416 m / 0.075kg = 5.55 m


We replace data in the formulas:


Freezing point depression:0° - Freezing point of solution= 1.86°C/m. 5.55 m


Freezing point of solution = - (1.86°C/m . 5.55m) = - 10.3°C


Boiling point elevation:Boiling point of solution - 100°C = 0.51°C /m . 5.55 m


Boiling point of solution = 0.51°C /m . 5.55 m + 100°C = 102.8°C